OpenStudy (anonymous):

A paint mixer wants to mix paint that is 15% gloss with paint that is 30% gloss to make 6 gallons that is 25% gloss. How many gallons of each paint should the paint mixer mix together?

hero (hero):

I just responded to a similar question earlier. x + y = 6 .15x + .30y = 6(.25)

OpenStudy (anonymous):

What do you do next ?

hero (hero):

This is systems of equations. You find x and y which gives you the amount of gallons of each paint

OpenStudy (anonymous):

Can you explain this step by step I'm having trouble trying to work this out.

hero (hero):

You've never done systems of equations before?

OpenStudy (anonymous):

I have but I can't remember how to start or anything . I'm doing poorly in algebra 1 right now :(

hero (hero):

Okay, but take notes... Using given equations: x + y = 6 .15x + .30y = 6(.25) Write each equation in terms of y as follows: y = 6 - x y = (1.5 - .15x)/.30 Next set expressions y = y, then cross multiply: .30(6 - x) = 1.5 - .15x Now solve for x: 1.8 - .30x = 1.5 = .15x .3 = .15x .3/.15 = x 2 = x 2 gallons of paint for the 15% gloss Now substitute x back into the original equation to solve for y: 2 + y = 6 y = 6 - 2 y = 4 4 gallons of paint for the 30% gloss

OpenStudy (anonymous):

Thanks

hero (hero):

I prefer if you think me with medals :D